Evaluate ſxsin ( x ) ſxsin (6)
Answer #1

(xsin (2) dx het u=2-0 du da 6 6 da=bdu 2 x=bu 36 usinu du S (6u) Sin (W) (6 du) =) * 36 usinudu Integrating by parts we get= -36 x Cos (2) + 36sin (2) te 6 - 6xlos (2) + 36 sin ( 2 ) + osin (2) 6x cos(x) + c xsin ydx= 36 6